# CleanSurvival: Automated data preprocessing for time-to-event models using reinforcement learning

### v1

@@ Abstract
Data preprocessing is a critical yet frequently neglected aspect of machine learning, often paid little attention despite its potentially significant impact on model performance. While automated machine learning pipelines are starting to recognize and integrate data preprocessing into their solutions for classification and regression tasks, this integration is lacking for more specialized tasks like survival or time-to-event models. As a result, survival analysis not only faces the general challenges of data preprocessing but also suffers from the lack of tailored, automated solutions in this area.
To address this gap, this paper presents CleanSurvival, a reinforcement-learning-based solution for optimizing preprocessing pipelines, extended specifically for survival analysis. The framework can handle continuous and categorical variables, using $Q$-learning to select which combination of data imputation, outlier detection and feature extraction techniques achieves optimal performance for a Cox, random forest, neural network or user-supplied time-to-event model. The package is available on GitHub: https://github.com/datasciapps/CleanSurvival.
Experimental benchmarks on real-world datasets show that the $Q$-learning-based data preprocessing results in superior predictive performance to standard approaches, finding such a model up to 10 times faster than undirected random grid search. Furthermore, a simulation study demonstrates the effectiveness in different types and levels of missingness and noise in the data.
@@ 1 Introduction
Survival analysis poses unique challenges due to the presence of censored data, where the event of interest has not yet been observed. However, it is also overlooked in the context of automated machine learning (AutoML) pipelines, which aim to streamline the ML development process by automating tasks such as algorithm selection, hyperparameter tuning and model evaluation—and increasingly incorporate data preparation as part of their pipelines. In this paper, we introduce CleanSurvival, an automated data preprocessing framework tailored for survival analysis.
@@
The article is organized as follows. Section 2 provides an overview of data preprocessing, survival analysis, AutoML and $Q$-learning. Section 3 reviews existing approaches to automated data preprocessing and AutoML frameworks. Section 4 describes the architecture of CleanSurvival and its features. Section 5 presents the results of experimental evaluation of the framework on real-world datasets. Finally, Section 6 discusses the results and outlines future directions for research.
@@ 2 Background
Data preprocessing involves cleaning, transforming and organizing raw data into a suitable format for analysis, and is a important step in the ML pipeline. Sub-tasks of data preprocessing include imputation or removal of missing values, detection and handling of outliers, variable selection and feature extraction and data transformations. These steps can have a profound downstream impact on classification performance [15] and model explanations [30].
@@ 2.1 Survival analysis
Survival analysis models, such as the Cox proportional hazards model, Kaplan–Meier estimator and accelerated failure time models, are widely used in practice. These models estimate the hazard function, survival function, or survival probabilities over time, providing valuable insights into the relationship between covariates and survival outcomes.
Concordance indices like the C-index are widely used in evaluation of survival models due to their simplicity and ease of interpretation [16]. The C-index evaluates the model’s ability to correctly rank individuals based on their risk of experiencing the event. A higher C-index indicates better discriminatory power.
However, concordance indices have significant limitations, due to their poor calibration and failure to consider the distribution of survival times as well as their ranks. To measure how well survival probabilities align with observed event times, calibration metrics like Houwelingen’s $\alpha$ or D-calibration can be used [35] , and should be compared against a baseline model, such as the Kaplan–Meier estimator. Weighted integrated survival log loss or integrated Graf score are recommended scoring rules [31].
@@
for a set of time points $\{t_{1},t_{2},\ldots,t_{k}\}$, where $\Delta t_{j}$ represents the difference between time points, $n$ is the number of individuals and $\text{Observed}(t,i)$ is an indicator function, equal to $1$ if the individual is known to have survived beyond $t$ and 00 otherwise.
@@ 2.2 Automated machine learning
However, data preprocessing remains an important analysis step that typically falls outside the AutoML pipeline [23]. Data preparation steps including cleaning, normalization and feature engineering are critical for the success of ML models but can be highly problem-specific [12]. Automating these tasks while maintaining flexibility for diverse datasets remains a significant hurdle [27, 19]. Mahdavi et al. [18] highlighted the potential of AI to solve data quality problems through data profiling and learning from past cleaning attempts [see also 17]. A holistic approach integrates the cleaning process with downstream tasks so that the cleaning is optimized for predictive performance [20]; indeed when integrated into an AutoML framework some cleaning steps may be more important than others [21].
@@
Survival analysis in particular faces particular challenges, partly due to the relative lack of support for such models in the first place (versus classification or regression), as well as unique difficulties of handling censored time-to-event data [36], which are typically not addressed in conventional AutoML frameworks and do not feature in AutoML surveys [e.g. 1]. Seamlessly integrating these domain-specific preprocessing steps with the downstream tasks of model optimization and evaluation is a complex, underexplored area.
@@
Prominent AutoML pipelines include Auto-WEKA, an early AutoML system that uses Bayesian optimization to search for the best combination of preprocessing steps and machine learning algorithms [34]; TPOT, a tree-based pipeline optimization tool that uses genetic programming to evolve pipelines of data cleaning and machine learning operations [22]; and auto-sklearn, an extension of Auto-WEKA that incorporates more recent advancements in machine learning and hyperparameter optimization while offering a familiar interface based on the Python package scikit-learn [8].
These AutoML pipelines have demonstrated promising results in various domains, including image classification, natural language processing and tabular data analysis. However, they often focus on general machine learning tasks and may not be specifically tailored to the challenges of survival analysis, particularly in the presence of missing data.
@@
Salhi et al. [27] presented a recent survey of data preprocessing using AutoML (though survival analysis is not mentioned). In their review, they highlight the relative capabilities of AutoML platforms: in many cases the data processing support is relatively basic. The authors indicate that all 11 tools reviewed support missing value imputation, but also state: ‘auto-sklearn cannot handle missing values’ and this must be done manually by the user—a contradiction. In this case, claims of support for data processing may actually be based on those of the underlying, non-automated ML framework (i.e. scikit-learn). Mumuni and Mumuni [19] also surveyed automated data processing for deep learning applications, highlighting in more detail the extent to which data processing steps are integral components of the automated pipeline. They similarly note the lack of early support from autosklearn.
@@ 2.3 $Q$-learning
3 Related work
Table 1 compares the features of various AutoML solutions; as highlighted in Section 2, few offer integrated support for survival analysis and data preprocessing is not always within the optimization loop.
Table 1: Comparison of AutoML frameworks and their native support for survival analysis and dynamic data preprocessing
Framework Method Preprocessing Survival
Amazon SageMaker Autopilot Bayesian optimization and ensembles Yes No
auto-keras Neural architecture search Yes No
auto-sklearn Bayesian optimization Limited No
AutoGluon Stack ensembling Yes No
Azure AutoML Bayesian optimization and meta-learning Yes No
BigML Decision tree-based optimization Yes No
DataRobot Proprietary ensemble and optimization Yes Limited
FLAML Cost-aware Bayesian optimization No No
Google AutoML Tables Neural architecture search Yes No
H2O AutoML Random search and stacked ensembles Yes No
MLflow Manual configuration Yes No
MLJAR Random search and stacked ensembles Yes No
PyCaret Iterative search with pipeline tuning Yes Limited
TPOT Genetic programming No No
Of frameworks offering automated data cleaning, DataRobot is proprietary, commercial platform and only appears to offer time-to-event modelling via a ‘hack’ of converting the task to a classification problem via discretization [see e.g. 5]. Meanwhile, H2O.ai can run Cox proportional hazards models as a fixed model, but not via its AutoML interface.
However, some dedicated data cleaning solutions have been proposed. Bilal et al. [3] proposed Auto-Prep, an interactive Python-based tool that recommends data cleaning methods to the user based on application of candidate techniques and subsequent evaluation using simple classifiers or regression models. In a review of data preprocessing in AutoML (Section C) the authors highlight the capabilities, or lack thereof, of AutoML tools to perform data preprocessing and feature engineering without manual human intervention. Another Python package, Atlantic [28], automates preprocessing steps including feature engineering and missing value imputation for supervised learning tasks. The framework identifies the best combination of steps based on evaluation using tree-based model ensembles.
Berti-Equille [2] developed Learn2Clean, a tool offering an innovative approach to data preprocessing. It leverages $Q$-Learning, a reinforcement learning technique, to dynamically select the optimal sequence of preprocessing tasks for a given dataset and ML model. This optimization aims to maximize the quality of the ML model’s results. Learn2Clean implements automated data preprocessing for regression, classification and clustering tasks, using $Q$-learning to optimize respective evaluation criteria: mean squared error, accuracy and silhouette index. However, Learn2Clean limitations include lack of built-in support for survival analysis, categorical data types, flexible hyperparameter tuning and custom reward functions. It also has a complex dependency structure, which can make initial setup challenging for end-users.
@@ 4.1.1 Missing values
  • Multiple imputation using chained equation [MICE, 4] provides a more robust approach by generating multiple suitable replacements for each missing value, creating several complete datasets for analysis. This method utilizes an iterative imputer, which starts with an initial guess and refines the estimates until convergence.
@@ 4.1.2 Outliers
To ensure the reliability and validity of survival analysis, robust outlier detection methods are essential.
@@ 4.3 Working modes
In this mode, users can specify the desired number of random experiments. The tool generates random preprocessing pipelines and evaluates their performance, providing insights into the impact of different preprocessing choices. This mdoe can serve as a baseline for comparison with the optimized pipeline.
@@ 5.1 Experimental setup
All methods were implemented in Python. Source code and documentation are available at https://github.com/datasciapps/CleanSurvival. Details of the datasets used are provided in Table 2. We demonstrate the approach using the rotterdam dataset, derived from the Rotterdam Study, a large prospective cohort study in the Netherlands, and flchain, a study of the relationship between serum free light chain (a type of blood measurement) and mortality.
@@
Flchain 7874 11 UCI Machine Learning Repository
@@
    complete case analysis (CCA), effectively ignoring the problem of missingness and outliers
@@
    random selection of imputation methods, simulating a grid search over possible analysis choices
@@
Hyperparameters that are not part of the CleanSurvival search space were left at default values, with architectural choices for the deep learning framework as follows: the DeepHit architecture consists of a shared sub-network with 2 hidden layers, each containing 100 neurons, using ReLU as the activation function. This is followed by cause-specific sub-networks, each containing 2 hidden layers with 50 neurons per layer. Dropout regularization was applied to all layers with a keep probability of 0.8 to mitigate overfitting. The model was trained using the Adam optimizer with a learning rate of 0.001 for 10,000 iterations. In practice, these options are customizable by the user or could be incorporated into the reinforcement learning action space as additional modules.
Since the Rotterdam dataset does not contain missing values, three variations of the dataset were used each representing a different missingness strategy (MCAR, MAR, MNAR) which were generated using the Jenga framework [29].
@@ 5.2 Results
(a) MCAR
(b) MAR
(c) MNAR
(d) Flchain
@@
These findings highlight the potential of CleanSurvival as a valuable tool for accelerating and automating the data preprocessing stage of survival analysis. By efficiently identifying optimal pipelines, not only does it save time and resources but it also enables researchers to focus on the core aspects of their analysis, ultimately leading to more robust and reliable results.

### v2

@@ Abstract
Background: Data preprocessing is a critical yet frequently neglected aspect of machine learning, often paid little attention despite its potentially significant impact on model performance. While automated machine learning pipelines are starting to recognize and integrate data preprocessing into their solutions for classification and regression tasks, this integration is lacking for more specialized tasks like survival or time-to-event models. As a result, survival analysis not only faces the general challenges of data preprocessing but also suffers from the lack of tailored, automated solutions in this area.
Method: To address this gap, this paper presents CleanSurvival, a reinforcement-learning-based solution for optimizing preprocessing pipelines, extended specifically for survival analysis. The framework can handle continuous and categorical variables. It builds upon Learn2Clean’s $Q$-learning to select which combination of data imputation, outlier detection and feature extraction techniques achieves optimal performance for a Cox, random forest, neural network or user-supplied time-to-event model. The package is available on GitHub: https://github.com/phoenix0401/CleanSurvival.
Results: Experimental benchmarks on real-world datasets show that the $Q$-learning-based data preprocessing results in superior predictive performance to standard approaches, finding such a model up to 10 times faster than undirected random grid search. Furthermore, a simulation study demonstrates the effectiveness in different types and levels of missingness and noise in the data.
Conclusion: With an increase in the use of machine learning, it becomes important to generalise AutoML pipelines to a variety of models now present, including survival analysis. Tools like CleanSurvival, which integrate preprocessing for survival analysis, can make survival studies easier and quicker to perform, as well as make the results more robust.
@@ 1 Introduction
Survival analysis poses unique challenges due to the presence of censored data, where the event of interest has not yet been observed. However, it is also overlooked in the context of automated machine learning (AutoML) pipelines, which aim to streamline the ML development process by automating tasks such as algorithm selection, hyperparameter tuning and model evaluation—and increasingly incorporate data preparation as part of their pipelines. There exist several tools focusing on preprocessing for classification and regression models [27]. In this paper, we introduce CleanSurvival, an automated data preprocessing framework tailored specifically for survival analysis.
@@
The article is organized as follows. Section 2 provides an overview of data preprocessing, survival analysis, AutoML tools and $Q$-learning. Section 3 describes the architecture of CleanSurvival and its features. Section 4 presents the results of experimental evaluation of the framework on real-world datasets and Section 5 discusses CleanSurvival in comparison to other upcoming tools in this field. Finally, Section 6 discusses the results and outlines future directions for research.
@@ 2 Background
Data preprocessing involves cleaning, transforming and organizing raw data into a suitable format for analysis, and is a important step in the ML pipeline. Sub-tasks of data preprocessing include imputation or removal of missing values, detection and handling of outliers, variable selection, feature extraction and data transformations. These steps can have a profound downstream impact on classification performance [22] and model explanations [40].
@@ 2.1 Survival analysis
Survival analysis models, such as the Cox proportional hazards model, Kaplan–Meier estimator and accelerated failure time models, are widely used in practice [50, 41]. These models estimate the hazard function, survival function, or survival probabilities over time, providing valuable insights into the relationship between covariates and survival outcomes.
Concordance indices like the C-index are widely used in the evaluation of survival models due to their simplicity and ease of interpretation [23]. The C-index evaluates the model’s ability to correctly rank individuals based on their risk of experiencing the event. A higher C-index indicates better discriminatory power.
However, concordance indices have significant limitations, due to their poor calibration and failure to consider the distribution of survival times as well as their ranks [16]. The Brier Score is another metric in survival analysis which calculates the accuracies of the predicted survival probabilities at a given time point [42]. Other calibration metrics like Houwelingen’s $\alpha$ or D-calibration can also be used to measure how well survival probabilities align with observed event times [49]. They could be compared against a baseline model such as the Kaplan–Meier estimator. Weighted integrated survival log loss or integrated Graf score are recommended scoring rules [43].
@@
for a set of time points $\{t_{1},t_{2},\ldots,t_{k}\}$, where $\Delta t_{j}$ represents the difference between time points, $n$ is the number of individuals and $\text{Observed}(t,i)$ is an indicator function, equal to $1$ if the individual is known to have survived beyond $t$ and 0 otherwise.
@@ 2.2 Automated machine learning
However, data preprocessing remains an important analysis step that typically falls outside the AutoML pipeline [32]. Data preparation steps including cleaning, normalization and feature engineering are critical for the success of ML models but can be highly problem-specific [19]. Automating these tasks while maintaining flexibility for diverse datasets remains a significant hurdle [37, 27]. Mahdavi et al. [25] highlighted the potential of AI to solve data quality problems through data profiling and learning from past cleaning attempts ([24]. A holistic approach integrates the cleaning process with downstream tasks so that the cleaning is optimized for predictive performance [29]; indeed when integrated into an AutoML framework, some cleaning steps may be more important than others [30].
@@
Survival analysis in particular faces particular challenges, partly due to the relative lack of support for such models in the first place (versus classification or regression), as well as unique difficulties of handling censored time-to-event data [50], which are typically not addressed in conventional AutoML frameworks and do not feature in AutoML surveys [2]. Seamlessly integrating these domain-specific preprocessing steps with the downstream tasks of model optimization and evaluation is a complex, underexplored area.
@@
Prominent AutoML pipelines include Auto-WEKA, an early AutoML system that uses Bayesian optimization to search for the best combination of preprocessing steps and machine learning algorithms [48]. It covers a range of models but does not cover crucial survival models like Cox or survival metrics [18]. TPOT, a tree-based pipeline optimization tool that uses genetic programming to evolve pipelines of data cleaning and machine learning operations, however it performs simple imputation and does not perform censoring natively which means it cannot work on survival analysis [31]. Similarly, auto-sklearn, an extension of Auto-WEKA that incorporates more recent advancements in machine learning and hyperparameter optimization while offering a familiar interface based on the Python package scikit-learn cannot perform preprocessing for survival models and metrics [13].
@@
[37] presented a recent survey of data preprocessing using AutoML (though survival analysis is not mentioned). In their review, they highlight the relative capabilities of AutoML platforms: in many cases the data processing support is relatively basic. The authors indicate that all 11 tools reviewed support missing value imputation, but also state: ‘auto-sklearn cannot handle missing values’ and this must be done manually by the user—a contradiction. In this case, claims of support for data processing may actually be based on those of the underlying, non-automated ML framework (i.e. scikit-learn). [27] also surveyed automated data processing for deep learning applications, highlighting in more detail the extent to which data processing steps are integral components of the automated pipeline. They similarly note the lack of early support from autosklearn.
Table 1: Comparison of AutoML frameworks with respect to data preprocessing and survival analysis capabilities
Framework License Missing values Categorical Imbalance Survival
auto-sklearn Open-source Requires NaN removal Yes Yes External libraries
TPOT Open-source Yes No No No
Google AutoML Tables Proprietary Yes Yes Yes No
H2O AutoML Open-source Yes Yes Yes Limited
PyCaret Open-source Yes Yes Yes Limited
AutoKeras Open-source Yes Yes Limited Custom models
Azure AutoML Proprietary Yes Yes Yes External tools
BigML Commercial Yes Yes Limited No
MLflow Open-source No No No No
FLAML Open-source No No Yes No
MLJAR Open-source Yes Yes Limited No
DataRobot Proprietary Yes Yes Yes Limited
Amazon SageMaker Autopilot Proprietary Yes Yes Yes No
AutoGluon Open-source Yes Yes Limited External libraries
Table 1 compares the features of various AutoML solutions. Of frameworks offering automated data cleaning, DataRobot is proprietary, commercial platform and only appears to offer time-to-event modelling via a ‘hack’ of converting the task to a classification problem via discretization (see e.g. [9]). Google AutoML Tables, Amazon SageMaker Autopilot and Azure ML can perform different preprocessing steps but they are not always tailored to survival analysis and these are all proprietary tools [28]. Big ML, that performs decision tree-based optimization, performs preprocessing but no survival analysis [12]. MLJAR and AutoGluon both use stack ensembling, perform most of the preprocessing steps needed but do not suppost survival analysis [26, 46]. Meanwhile, PyCaret can perform preprocessing and survival analysis as well but only to an extent. It can do mortality predictions but needs external tools for further survival analysis [35]. Likewise, H2O.ai can run Cox proportional hazards models [8] as a fixed model, but not via its AutoML interface as it requies advanced platforms for the same.
However, some dedicated data cleaning solutions have been proposed. [4] proposed Auto-Prep, an interactive Python-based tool that recommends data cleaning methods to the user based on application of candidate techniques and subsequent evaluation using simple classifiers or regression models. In a review of data preprocessing in AutoML (Section C) the authors highlight the capabilities, or lack thereof, of AutoML tools to perform data preprocessing and feature engineering without manual human intervention. Another Python package, Atlantic [38], automates preprocessing steps including feature engineering and missing value imputation for supervised learning tasks. The framework identifies the best combination of steps based on evaluation using tree-based model ensembles.
@@ 2.3 $Q$-learning
Berti-Equille [3] developed Learn2Clean, a tool offering an innovative approach to data preprocessing. It leverages $Q$-Learning, a reinforcement learning technique, to dynamically select the optimal sequence of preprocessing tasks for a given dataset and ML model. This optimization aims to maximize the quality of the ML model’s results. Learn2Clean implements automated data preprocessing for regression, classification and clustering tasks, using $Q$-learning to optimize respective evaluation criteria: mean squared error, accuracy and silhouette index. However, Learn2Clean limitations include lack of built-in support for survival analysis, categorical data types, flexible hyperparameter tuning and custom reward functions. It also has a complex dependency structure, which can make initial setup challenging for end-users. We build upon Learn2Clean, by adapting the reinforcement learning framework to survival analysis.
@@ 3.1.1 Missing values
  • Multiple imputation using chained equation [5] provides a more robust approach by generating multiple suitable replacements for each missing value, creating several complete datasets for analysis. This method utilizes an iterative imputer, which starts with an initial guess and refines the estimates until convergence.
@@ 3.1.2 Outliers
To ensure the reliability and validity of survival analysis, robust outlier detection methods are essential [6].
@@ 3.3 Working modes
In this mode, users can specify the desired number of random experiments. The tool generates random preprocessing pipelines and evaluates their performance, providing insights into the impact of different preprocessing choices. This mode can serve as a baseline for comparison with the optimized pipeline.
@@ 4.1 Experimental setup
All methods were implemented in Python. Source code and documentation are available at https://github.com/phoenix0401/CleanSurvival. Details of the datasets used are provided in Table 2. We demonstrate the approach using the rotterdam dataset, derived from the Rotterdam Study, a large prospective cohort study in the Netherlands, and flchain, a study of the relationship between serum free light chain (a type of blood measurement) and mortality.
@@
Flchain 7874 11 UCI Repository
@@
    Complete case analysis (CCA): effectively ignoring the problem of missingness and outliers
@@
    Random selection of imputation methods: simulating a grid search over possible analysis choices
@@
Hyperparameters that are not part of the CleanSurvival search space were left at default values. The architectural choices for the deep learning framework were as follows: the DeepHit architecture consists of a shared sub-network with 2 hidden layers, each containing 100 neurons, using ReLU as the activation function. This is followed by cause-specific sub-networks, each containing 2 hidden layers with 50 neurons per layer. Dropout regularization was applied to all layers with a keep probability of 0.8 to mitigate overfitting. The model was trained using the Adam optimizer with a learning rate of 0.001 for 10,000 iterations. In practice, these options are customizable by the user or could be incorporated into the reinforcement learning action space as additional modules.
Since the Rotterdam dataset does not contain missing values, three variations of the dataset were used each representing a different missingness strategy (MCAR, MAR, MNAR) which were generated using the Jenga framework [39]. After generating pipelines using CleanSurvival and Random, we compute Brier scores across timepoints (n=50) for the Rotterdam dataset, with three mechanisms and varying levels of missingness.
@@ 4.2 Results
The C-index values in Table 3 show that CleanSurvival leads to better performance of the models across datasets with varying degrees of missingness. It has a higher C-index compared to CCA and Mean imputation. With Rotterdam datasets having high C-index and Flchain, this could be due to the different mechanisms of simulated missingness versus the natively-present missing values in FlChain.
Refer to caption Refer to caption Refer to caption
(a) MCAR (b) MAR (c) MNAR
Figure 2: CleanSurvival vs Random
Figure 2 shows the Brier scores over time for various missingness percentages as well as for the three mechanisms used to simulate missingness in the Rotterdam dataset. With timepoints chosen to be 50, it can be seen that overall CleanSurvival generated more stable Brier scores compared to the baseline of randomly generated pipelines. For the MCAR mechanism of missingness, across all levels of missingness, the Brier scores are stable and end at a lower score compared to Random. This is visible for both types of models - COX and RSF. For MAR and MNAR, apart from one missingness level (40 for the former and 30 for the latter), other line plots are all stable for the COX and RSF models when pipelines were preprocessed with CleanSurvival. These plots show that CleanSurvival can help improve robustness and consistency across different models and runs.
Refer to caption Refer to caption Refer to caption Refer to caption
(a) MCAR (b) MAR (c) MNAR (d) Flchain
@@
These findings highlight the potential of CleanSurvival as a valuable tool for accelerating and automating the data preprocessing stage of survival analysis. By efficiently identifying optimal pipelines, not only does it save time and resources, but it also enables researchers to focus on the core aspects of their analysis, ultimately leading to more robust and reliable results.
5 Discussion
With most of the AutoML preprocessing tools focusing on classification and regression models, we present CleanSurvival as a tool which builds upon Learn2Clean to tailor preprocessing for survival analysis using a $Q$-learning-based framework and reward structures. New tools like AutoPrognosis [1] have developed automated tools using Bayesian Optimization that include survival analysis, but there is less scope for customisation of the pipelines. The widely used library scikit-survival integrates scikit-learn preprocessing with survival models but it requires manual pipeline building [7]. The R library ‘mlr3proba’ provides a machine learning interface for survival analysis and preprocessing, however missingness is still dealt with externally [44]. CleanSurvival can be performed on various survival models, including neural networks. It can handle different preprocessing steps required for survival analysis data. The user can also customise pipelines based on their requirement. This makes CleanSurvival an important tool in the emerging field of automation of machine learning and preprocessing pipelines.